# An Economic Solution to Copyright Challenges of Generative AI

### v3

@@ 1 Introduction
Efforts to mitigate the tension between owners of copyright in the training data and AI developers have emerged, mostly involving modifications to generative model training or inference to reduce the likelihood of generating infringing outputs [35, 4, 33]. However, these modifications may compromise model performance due to either the exclusion of high-quality, copyrighted training data from training or restrictions on content generation [19]. The complexity and ambiguity of copyright law add another layer of difficulty, blurring the line between infringing and non-infringing outputs. The resulting uncertainty could lead to a significant waste of resources on both sides while these issues are debated in courts [32].
@@ 2 The Shapley Royalty Share Framework
The utility offers a way to measure the extent to which the data sources from $S$ are responsible for generating the content. It is small if the counterfactual model is unlikely to generate the same content as the deployed model, and vice versa.
@@
The utility (2.1) or (2.2) can be interpreted as the total compensation all members of $S$ collectively deserve for providing their data to train the generative AI model. The next step is to determine the payoff for each individual copyright owner, based on the utilities of all possible combinations of data sources. We propose using the Shapley value [34] for this task. The Shapley value is a solution concept in cooperative game theory that offers a principled approach to distributing gains depending on the utility of every combination of players as a coalition. It is the only payment rule satisfying several important economic properties (see the supplementary materials for details) [34, 29] and has gained popularity in data valuation for machine learning models [9, 16].
@@
Computational considerations
A main challenge in applying the framework of SRS lies in its substantial computational cost. The evaluation of the utility functions on different combinations of data sources requires retraining the model multiple times. In some applications where the number of copyright owners is small, the computational challenge might not be as severe as it seems. Indeed, we envision that this contract-based framework works best when the entire copyrighted data is partitioned among a handful of copyright owners so that each source has enough data to impact the training outcome. If the data source is very small in size, the royalty share of the owner would be mostly insignificant and, worse, noisy due to the stochastic nature of training AI models [36].
To alleviate such computational burdens, two approaches can be applied here. The first is to use the Monte Carlo method to approximate the Shapley value [16, 15, 26, 38, 3, 25, 23, 37]. This technique is specially tailored to the case of a large number of copyright owners in the coalition. The second approach is to train a model by fine-tuning it from another model that is trained on a smaller subset of data. Hence, one can approximate models trained on different subsets of data sources by training the model with only one pass through the entire training data. Specifically, for a randomly sampled permutation of copyright owners, we can first train on the first copyright owner, then the second, and all the way up to the last copyright owner. This technique can be used together with the famous permutation sampling estimator for the Shapley value [24].
In practice, a commercial AI model could undergo millions of transactions on a daily basis. It suffices to estimate the aggregated payoffs each copyright owner deserves instead of calculating the payoff as specified in (2.4) for each AI-generated content. To save computational cost, we can evaluate the SRS for only a small fraction of all transactions and scale back to obtain estimates of the revenue distributions from all transactions; see a detailed discussion in the supplementary materials.
@@
3 Results
We assessed the proposed framework’s effectiveness in distributing royalties for AI-generated content using experiments, with a focus on creative art and logos in the image domain. Our evaluation utilized publicly available datasets: WikiArt [31] and FlickrLogo-27 [17], with detailed dataset and training algorithm settings provided in the supplementary materials.
Evaluation protocol
For the WikiArt dataset, we selected four disjoint subsets of paintings from four renowned artists. A model, initially trained on a broader set of training images (excluding those belonging to the four artists), served as the base model. The SRS is computed by further fine-tuning the base model on various combinations of the four painting sets belonging to the selected artists. Similarly, for the FlickrLogo-27 dataset, we selected four disjoint subsets of logo designs from four brands, and computed the SRS using a base model trained on logo images from other brands. Our goal was to assess whether the SRS can reflect each copyright owner’s contribution to the generation of images.
@@ 3 Results
Figure 2 shows the computed SRS for different kinds of $x^{(\text{gen})}$’s that are either the original or the AI-generated painting that is in the style of different artists. The results indicate that the SRS has the highest values when the $x^{(\text{gen})}$’s closely resembles the training data source in style. This relationship underscores the SRS framework’s ability to accurately attribute contributions to the creation of AI-generated images.
@@
In Figure 3, we explored the SRS distribution for prompts requesting content generation from multiple data sources. Notably, for the WikiArt dataset, prompts asked the generative model to blend styles from multiple artists. The SRS effectively recognized and rewarded the contributions of data sources integrated into the generated artworks, showcasing the framework’s capability to discern and value diverse data source inputs to generate content.
@@
We further explored the SRS framework’s response to prompts requesting content generation from non-copyrighted data sources, as shown in Figure 4. In these scenarios, the SRS distribution was observed to be nearly uniform across all copyright owners. This outcome aligns with expectations, as the generated content lacks direct ties to any of the copyrighted data sources. This uniformity demonstrates the SRS framework’s ability to avoid disproportionate revenue distribution.
@@ 4 Related Work
Recent efforts in machine learning have primarily focused on minimizing the likelihood of creating copyright-infringing content by generative AI models. One approach involves training an auxiliary generative model on non-copyrighted data and utilizing rejection sampling to reduce the likelihood of reproducing copyrighted material [35]. However, this method is susceptible to adversarial attacks [22]. Alternatively, [4] suggests modifying generative models’ training objectives to avoid generating outputs that closely resemble copyrighted data. Yet another technique focuses on protecting unique artistic styles by incorporating adversarial perturbations into copyrighted images for model fine-tuning [33].
The Shapley value has been suggested as a means to fairly distribute revenue in traditional sectors such as royalty agreements between music copyright holders and radio broadcasters [39]. The Shapley value has been used for data valuation where the utility function is the prediction accuracy of the machine learning model [9, 16]. This differs from our SRS framework, which uses the log-likelihood as the utility since there is no such thing as prediction accuracy for generative models.
Other data valuation techniques have been developed for generative models. A simple approach utilizes similarity scores between training data and generated content as a valuation metric [40]. Another commonly used approach is the leave-one-out (LOO) score or its approximations. For example, [8] extends the TRAK framework [27] to generative models, and [41] further introduced empirical approaches to improving the performance of [8]. Notably, [6] proposed a revenue-sharing mechanism for AI-generated music based on TRAK, which is closely related to our work. However, the LOO scores neglect the high-order training data interactions, which may result in undesirable attribution scores (see Appendix B for detailed discussion).
@@ A.1 Efficient SRS Estimation for High-volume Transactions
Considering the potential for millions of daily uses and transactions of a commercial AI model, computing the Shapley Royalty Share (SRS) for each individual transaction may not be feasible due to the high computational cost. To address this challenge, a practical solution is to estimate the average SRS for each copyright owner based on a subset of transactions. For any given transaction involving AI-generated content $x^{(\text{gen})}$, let $p_{x^{(\text{gen})}}$ represent the payment by the user. The share of this payment received by copyright owner $i$ is determined by:
@@ Appendix C Experimental Settings
The generative models used in our experiment follow from the recent advancements in high-resolution image synthesis utilizing latent diffusion models [28]. [^3] Part of the codebase is adapted from https://github.com/artem-gorodetskii/WikiArt-Latent-Diffusion and https://github.com/VSehwag/minimal-diffusion. All images were cropped and resized to $512\times 512$ resolution. As training a new generative model from scratch would be prohibitively costly, we use LoRA [14], an efficient fine-tuning method that enables us to scale up the models being used in a tractable manner.444 The implementation is adapted from https://huggingface.co/blog/lora. The model used for fine-tuning is Stable Diffusion V1-4555https://huggingface.co/CompVis/stable-diffusion-v1-4., which is a latent image diffusion model trained on LAION2B-en [^6] https://huggingface.co/datasets/laion/laion2B-en.. For the WikiArt dataset, we train each model on painting images from the same artist, with the text prompt “A painting in the style of [artist name].” Similarly, for the FlickrLogo-27 dataset, we train each model on logo images from the same artist, with the text prompt “A logo by [company name].” For each dataset, the model is fine-tuned with an initial learning rate $10^{-4}$, 10 epochs, and batch size $4$. We train the diffusion models on A100 GPU cluster. For all the SRS results in the maintext, they are being averaged over 20 $x^{(\text{gen})}$’s that are randomly selected/generated according to the specifications in the maintext.
@@
where $\mu_{\theta}(x_{t},t)$ and $\Sigma_{\theta}(x_{t},t)$ are the mean and covariance of the Gaussian distribution at step $t$, learned by the model with parameters $\theta$777 We note that for text-to-image diffusion models, $\mu_{\theta}(\cdot)$ and $\Sigma_{\theta}(\cdot)$ will also depend on the input text encoding, but the core mathematical representation remains the same. The model is trained to minimize the difference between the noisy data and its prediction of the denoised data at each step. This can be formalized as minimizing a loss function, e.g., the mean squared error (MSE), between the original data and the reconstructed data:

### v4

@@ 1 Introduction
Efforts to mitigate the tension between owners of copyright in the training data and AI developers have emerged, mostly involving modifications to generative model training or inference to reduce the likelihood of generating infringing outputs [37, 5, 35, 15, 4]. However, these modifications may compromise model performance due to either the exclusion of high-quality, copyrighted training data from training or restrictions on content generation [22]. The complexity and ambiguity of copyright law add another layer of difficulty, blurring the line between infringing and non-infringing outputs. The resulting uncertainty could lead to a significant waste of resources on both sides while these issues are debated in courts [34].
@@ 2 The Shapley Royalty Share Framework
This likelihood function serves as a natural utility measure for generative models, inspired by the fundamental principles of maximum likelihood estimation that underpin the training of many such models. The utility offers a way to measure the extent to which the data sources from $S$ are responsible for generating the content. It is small if the counterfactual model is unlikely to generate the same content as the deployed model, and vice versa. If the learning algorithm $\mathcal{A}$ is stochastic, we account for this randomness by taking the expectation over the algorithm. In such cases, the utility is defined as $v(S;x^{(\text{gen})}):=\operatorname{\mathbb{E}}_{\mathcal{A}}[\log p_{S}(x^{(\text{gen})})]$.
@@
The utility (2.1) or (2.2) can be interpreted as the total compensation all members of $S$ collectively deserve for providing their data to train the generative AI model. The next step is to determine the payoff for each individual copyright owner, based on the utilities of all possible combinations of data sources. We propose using the Shapley value [36] for this task. The Shapley value is a solution concept in cooperative game theory that offers a principled approach to distributing gains depending on the utility of every combination of players as a coalition. It is the only payoff rule that satisfies several crucial economic properties [36, 31], including efficiency, symmetry, dummy, and linearity. These properties are essential for ensuring fair revenue distribution among copyright owners (see supplementary materials for details).
@@
Computational considerations.
A main challenge in implementing SRS is the need to retrain the model multiple times to evaluate utility functions for different data source combinations. However, in applications with a few copyright owners, this challenge is much less pronounced. We envision this contract-based framework working best when copyrighted data is partitioned among a handful of owners, ensuring each source significantly impacts the training outcome. For very small data sources, the royalty share would be minimal and potentially noisy due to the stochastic nature of AI model training [38].
Additionally, for practical implementation in commercial AI models handling millions of daily transactions, estimating aggregated payoffs for each copyright owner, rather than calculating individual payoffs for each AI-generated content, would suffice. Computational costs can be further reduced by evaluating SRS for a small fraction of transactions and scaling to estimate overall revenue distributions.
Our work’s primary contribution lies in the novel concept of using the Shapley value to address the AI copyright issue, rather than in tackling these computational challenges. The latter remains an active research area in machine learning, including Monte Carlo methods for approximating the Shapley value [19, 18, 28, 41, 3, 27, 26, 39], particularly useful for large coalitions of copyright owners. Another approach to alleviating computational burdens in calculating Shapley values is the In-Run Data Shapley method [40]. This method significantly reduces computational overhead by computing Shapley values during the training process, rather than requiring multiple retraining runs on different data subsets. Our SRS framework lays the conceptual foundation for using Shapley values in AI copyright disputes, while acknowledging that efficient computation of these values remains an important area for future research and practical implementation.
@@
Remark 2.1 (Adversarial data owners).
We assume the data owners are honest and follow the protocol of SRS. In practice, an owner might attempt to artificially inflate their contribution by including non-copyrighted images or even AI-generated images in their dataset. Investigating robust methods to detect and mitigate such adversarial behavior will be crucial for ensuring the long-term integrity and fairness of the royalty-sharing framework in AI-generated content. These potential adversarial settings are important future directions and are beyond the scope of the current work.
3 Experiment
We assessed our proposed framework’s effectiveness in distributing royalties for AI-generated content across multiple copyright-sensitive domains: creative art, logo design, and language generation. We used three publicly available datasets: (1) WikiArt [33] for creative art, (2) FlickrLogo-27 [20] for logo design, and (3) Pile [9] for language generation. Detailed dataset descriptions and training algorithm settings are provided in the supplementary materials.
3.1 Evaluation protocol
Our evaluation protocol was designed to simulate real-world scenarios where AI models are trained on diverse datasets with multiple copyright owners. For each domain, we pretrain a model on the non-copyrighted dataset, followed by fine-tuning on copyrighted dataset. We finetune on different combinations of the copyrighted datasets to compute the SRS and assess its effectiveness in attributing contributions from various sources.
Creative Art (WikiArt).
We selected four disjoint subsets of paintings from four renowned artists (Van Gogn, Monet, Picasso, Rembrandt) as the copyrighted datasets. Our base model was trained on all artworks in the WikiArt dataset, excluding those belonging to the four selected artists.
Logo Design (FlickrLogo-27).
Similarly, we selected four disjoint subsets of logo designs from four brands (Sprite, Starbucks, Vodafone, Google). The SRS was computed using a base model trained on logo images from other brands not included in the selected four.
Language generation (Pile).
For the task of language generation, we use the Pile dataset. [^3] https://huggingface.co/datasets/monology/pile-uncopyrighted. Specifically, we first pretrain a Pythia 410M model on Pile-CC, a subset of the Pile [9] consisting of a general web corpus. We simulate the scenario where each owner of copyrighted data possesses a specific domain within the Pile, including PubMed Abstracts, PubMed Central, Wikipedia, GitHub, and EuroParl. We then estimate the Shapley values by fine-tuning the model on different combinations of these domains.
3.2 Results
@@
Figure 2 shows the computed SRS for different kinds of $x^{(\text{gen})}$’s that are either the original or the AI-generated painting that is in the style of different artists. The results indicate that the SRS has the highest values when the $x^{(\text{gen})}$’s closely resembles the training data source in style. This relationship underscores the SRS framework’s ability to accurately attribute contributions to the creation of AI-generated images. Figure 3 further illustrates this concept in the context of language models by displaying the Shapley values for different copyrighted subdomains of the Pile dataset in relation to generating a corpus in PubMed Abstract format. As expected, the PubMed Abstract subdomain achieves the highest Shapley value for the language model, while PubMed Central also scores highly due to domain similarity. Wikipedia, being a more general domain, shows a smaller but positive contribution to the model’s output. In contrast, EuroParl, which contains a significant amount of foreign language content, exhibits a negative Shapley value for the language model, likely due to the substantial domain shift from the target medical literature domain. More results are available in the supplementary materials (Section C.2).
@@
Figure 3: Left: The corpus from PubMed Abstract we use as the query example for contribution evaluation. Right: The Shapley value for the copyrighted owners.
Figure 4: Left: The corpus from NIH ExPorter, a medical dataset we use as the query example for contribution evaluation. Right: The Shapley value for the copyrighted owners.
Figure 5: Left: The corpus from DeepMind Mathematics (DM Mathematics), a corpus that contains simple math questions which we use as the query example for contribution evaluation. Right: The Shapley value for the copyrighted owners.
@@
In Figure 6, we explored the SRS distribution for prompts requesting content generation from multiple data sources. Notably, for the WikiArt dataset, prompts asked the generative model to blend styles from multiple artists. The SRS effectively recognized and rewarded the contributions of data sources integrated into the generated artworks, showcasing the framework’s capability to discern and value diverse data source inputs to generate content. Figure 4 illustrates the Shapley values when evaluating a corpus from NIH Exporter, a medical dataset similar to “PubMed Abstracts” and “PubMed Central”. As anticipated, both “PubMed Abstracts” and “PubMed Central” achieved the highest value scores, reflecting their close domain similarity to the NIH Exporter dataset. This result further validates the SRS framework’s capability to accurately attribute relevance to closely related data sources in specialized domains.
@@
We further explored the SRS framework’s response to prompts requesting content generation from non-copyrighted data sources, as shown in Figure 7. In these scenarios, the SRS distribution was observed to be nearly uniform across all copyright owners. This outcome aligns with expectations, as the generated content lacks direct ties to any of the copyrighted data sources. This uniformity demonstrates the SRS framework’s ability to avoid disproportionate revenue distribution. Figure 5 presents the Shapley values for generating a corpus in DM Mathematics. Notably, all domains exhibit very low or negative Shapley values, with the exception of Wikipedia. This exception can be attributed to Wikipedia’s substantial repository of math-related content, which likely contributes positively to the generation of mathematical text. This result underscores the SRS framework’s sensitivity to domain-specific relevance, even within broader, general-purpose datasets.
@@ 4 Related Work
Recent efforts in machine learning have primarily focused on minimizing the likelihood of creating copyright-infringing content by generative AI models. One approach involves training an auxiliary generative model on non-copyrighted data and utilizing rejection sampling to reduce the likelihood of reproducing copyrighted material [37]. However, this method is susceptible to adversarial attacks [25]. Alternatively, [5] suggested modifying generative models’ training objectives to avoid generating outputs that closely resemble copyrighted data. Yet another technique focuses on protecting unique artistic styles by incorporating adversarial perturbations into copyrighted images for model fine-tuning [35]. More recently, techniques such as rewriting the input prompt [15] and modifying outputs to be more generic [4] have also been proposed.
The Shapley value has been suggested as a means to fairly distribute revenue in traditional sectors such as royalty agreements between music copyright holders and radio broadcasters [42]. More recently, it has been used for data valuation where the utility function is the prediction accuracy of the machine learning model [11, 19]. This differs from our SRS framework, which uses the log-likelihood as the utility since there is no such thing as prediction accuracy for generative models.
There are a few techniques that have been extended to the generative models. A simple approach utilizes similarity scores between training data and generated content as a valuation metric [43]. Another commonly used approach is the leave-one-out (LOO) score or its approximations. For example, [10] extends the TRAK framework [29] to generative models, and [44] further introduced empirical approaches to improving the performance of [10]. Notably, [7] proposed a revenue-sharing mechanism for AI-generated music based on TRAK, which is closely related to our work. However, the LOO scores neglect the high-order training data interactions, which may result in undesirable attribution scores (see Section B in the supplementary materials for detailed discussion).
@@ A.1 Efficient SRS Estimation for High-volume Transactions
Considering the potential for millions of daily uses and transactions of a commercial AI model, computing the SRS for each individual transaction may not be feasible due to the high computational cost. To address this challenge, a practical solution is to estimate the average SRS for each copyright owner based on a subset of transactions. For any given transaction involving AI-generated content $x^{(\text{gen})}$, let $p_{x^{(\text{gen})}}$ represent the payment by the user. The share of this payment received by copyright owner $i$ is determined by:
@@ Appendix C Experimental Settings
The generative models used in our experiment follow from the recent advancements in high-resolution image synthesis utilizing latent diffusion models [30]. [^4] Part of the codebase is adapted from https://github.com/artem-gorodetskii/WikiArt-Latent-Diffusion and https://github.com/VSehwag/minimal-diffusion. All images were cropped and resized to $512\times 512$ resolution. As training a new generative model from scratch would be prohibitively costly, we use LoRA [17], an efficient fine-tuning method that enables us to scale up the models being used in a tractable manner.555 The implementation is adapted from https://huggingface.co/blog/lora. The model used for fine-tuning is Stable Diffusion V1-4666https://huggingface.co/CompVis/stable-diffusion-v1-4., which is a latent image diffusion model trained on LAION2B-en [^7] https://huggingface.co/datasets/laion/laion2B-en.. For the WikiArt dataset, we train each model on painting images from the same artist, with the text prompt “A painting in the style of [artist name].” Similarly, for the FlickrLogo-27 dataset, we train each model on logo images from the same artist, with the text prompt “A logo by [company name].” For each dataset, the model is fine-tuned with an initial learning rate $10^{-4}$, 10 epochs, and batch size $4$. We train the diffusion models on A100 GPU cluster. For all the SRS results in the maintext, they are being averaged over 20 $x^{(\text{gen})}$’s that are randomly selected/generated according to the specifications in the maintext.
@@
where $\mu_{\theta}(x_{t},t)$ and $\Sigma_{\theta}(x_{t},t)$ are the mean and covariance of the Gaussian distribution at step $t$, learned by the model with parameters $\theta$888 We note that for text-to-image diffusion models, $\mu_{\theta}(\cdot)$ and $\Sigma_{\theta}(\cdot)$ will also depend on the input text encoding, but the core mathematical representation remains the same. The model is trained to minimize the difference between the noisy data and its prediction of the denoised data at each step. This can be formalized as minimizing a loss function, e.g., the mean squared error (MSE), between the original data and the reconstructed data:
@@ C.1 The SRS of AI developers
C.2 Additional results
Identifying relevant copyright owners
Figure 10 presents the valuation results for a validation corpus derived from EuroParl. In this case, we observe that the copyright owner of the EuroParl dataset achieves a notably high Shapley value, while all other copyright owners’ Shapley values are close to zero. This stark contrast can be attributed to the fact that the EuroParl corpus contains significant non-English content, whereas the other datasets are predominantly in English. Figure 11 illustrates the valuation results for a validation corpus sourced from StackExchange, which is known to contain a substantial amount of code-related data. In this scenario, we find that the copyright owner of the Github dataset attains the highest Shapley value. This outcome is logical given that the Github dataset also comprises a large volume of code-related content, aligning closely with the nature of the StackExchange corpus. These results further demonstrate the SRS framework’s ability to accurately identify and valuate the most relevant data sources for different types of generated content, even when dealing with specialized or multilingual datasets.
Copyright owners with multiple datasets
In practice, a copyright owner may possess data from multiple sources. To simulate this scenario, we present in Figure 12 a case where one copyright owner holds a mixed dataset comprising EuroParl, FreeLaw, and NIH ExPorter. We evaluate this setup using a validation corpus sourced from PubMed Abstracts, which is known to contain a substantial amount of medical-related data. The results reveal that this mixed-dataset owner achieves a significantly higher Shapley value compared to an owner possessing only the EuroParl dataset (as shown in Figure 3). This marked increase can be attributed to the inclusion of NIH ExPorter in the mixed dataset, which also contains a significant amount of medical data. This demonstrates that the Shapley value framework can effectively capture the increased relevance and contribution of a copyright owner when they possess diverse data sources, particularly when these sources align with the domain of the generated content.
Figure 10: Left: The corpus from EuroParl, a multilingual dataset that contains European languages which we use as the query example for contribution evaluation. Right: The Shapley value for the copyrighted owners.
Figure 11: Left: The corpus from Stack Exchange Data Dump, which contains an anonymized set of all user-contributed content on the Stack Exchange network. Right: The Shapley value for the copyrighted owners.
Figure 12: Left: The corpus from PubMed Abstract we use as the query example for contribution evaluation. Right: The Shapley value for the copyrighted owners, where the owner ‘Mixture’ contains an equal amount of tokens from EuroParl, FreeLaw and NIH ExPorter.